OpenStudy (anonymous):

Help please: Micah is taking a test. Here is his answer to one of the problems. 34 × 33 = 97 Find and correct his error.

OpenStudy (anonymous):

its suppose to be 3^4*3^3=9^7

OpenStudy (texaschic101):

x^4 * x^3 = x^(m + n) 3^4 * 3^3 = 3^7